OpenStudy (anonymous):

I need help with these three, Assume that y varies inversely as x. If y = -6 when x = -2, find y when x = 5. Use the Law of the Lever to solve the following. Mark sits 4.5 feet from the center of the seesaw to balance a friend who weighs 53 lbs and is sitting 6 feet from the center. How many pounds does Mark weigh? Round to the nearest hundredth. Answer 39.75 lb 70.67 lb 55.78 lb Use the Law of Levers to solve the problem. The number of chairs on a ski lift varies inversely as the distance between them. A ski lift can handle 32 chairs when they are 30 feet apart. If 40 chairs were used, how many feet should be left between each pair? Answer 24 ft 37.5 ft 32 ft

jimthompson5910 (jim_thompson5910):

Assume that y varies inversely as x. If y = -6 when x = -2, find y when x = 5. "y varies inversely as x" translates to y = k/x y = k/x -6 = k/(-2) ... plug in y = -6 and x = -2 -6*(-2) = k 12 = k k = 12 the equation is y = 12/x now plug in x = 5 and evaluate y = 12/x y = 12/5 y = 2.4